OpenStudy (anonymous):

Choose the value of the y determinant (Dy) in the following system. 2x + 3y = 4

OpenStudy (anonymous):

Where did the second equation go?

OpenStudy (anonymous):

Yeah thats what im going to ask too?

OpenStudy (anonymous):

5x – y = 8 sorry this is it

OpenStudy (anonymous):

Use multiplication of the second equation by 3 and then add the two equations. From there, it is easy.

OpenStudy (anonymous):

yeah ^he got it..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Wait, wasn't the question the determinant for y? 2x + 3y = 4 5x - y = 8 Replace the coefficients of the y's with the values to the tight of the equal sign. 2x + 4 5x - y Remove the variables and set up a matrix. \[\left[\begin{matrix}2 & 4 \\ 5 & -1\end{matrix}\right]\]

OpenStudy (anonymous):

You are welcome. Have a good day now.

OpenStudy (anonymous):

@Calcmathlete, I'm getting tired. Maybe I goofed. I don't understand the use of a determinant in this context. Care to explain?

OpenStudy (anonymous):

Well you would use Cramer's Rule. http://www.purplemath.com/modules/cramers.htm They showed a 3 x 3 matix, but it works for 2 x 2 too.

OpenStudy (anonymous):

Irrelevant and unnecessary, in my opinion. However, from a pedagogical viewpoint, I guess the question is asking for that setup. Question: Are they asking for the determinant which evaluates to \(y\) or are they asking for the determinant of the system? I assume the former. What do you think?

OpenStudy (anonymous):

I concur. For this situation it's easier to solve by elimination or even substitution, but I do find that in some situations, it's easier to work with determinants that to work with he first two processes.

OpenStudy (anonymous):

Determinants are the systematic way of solving the equations. That's all. They are a more generalized solving technique. They're also very, very straightforward. There was an elegant theory behind all of this. However, I never learned this aspect, unfortunately. As such, I cannot quite appreciate the beauty of this technique.